2. If P(AB) 0.8 and P(AnB)-0.6, what is the PAcnB)? A) 0.75 B) 0.20C1.33D) 0.15
Answer #1

2)

Solution :

P(A | B) = 0.8

P(A \cap B) = 0.6

P(B) = P(A \cap B) / P(A | B) = 0.6 / 0.8 = 0.75

P(Ac\cap B) = P(B) - P(A \cap B) = 0.75 - 0.6 = 0.15